# Hatching-Box: Monitoring the Rearing Process of Drosophila Using an Embedded Imaging and in-vial Detection System

### v1

@@ Abstract
Drosophila melanogaster is among the most frequently used model organism in biology and life science as well as for development of novel treatments in the medical domain. In these fields, an analysis of behavioral patterns of the flies throughout their whole life cycle is of significant importance. For experimental purposes, large numbers of different Drosophila mutants are cultivated in rearing vials and monitored for anomalies in their behavior or development, for example when exposed to stimuli or substances. From this, researchers are able to draw conclusions regarding the effect of the involved substances or treatments on humans, other animals or the environment. A major drawback of current state-of-the-art monitoring systems is the high amount of manual and time-consuming labor as most of the methods employed today require individualization of the flies or similar steps which prevents fully automated experiments. In this paper we propose the Hatching-Box, a novel imaging and analysis system to automatically monitor and quantify the developmental behavior of Drosophila in standard rearing vials and during regular rearing routines, rendering explicit experiments obsolete. This is achieved by combining custom tailored imaging hardware with dedicated detection and tracking algorithms, enabling the quantification of larvae, filled/empty pupae and flies over multiple days. Given the affordable and reproducible design of the Hatching-Box in combination with our generic client/server-based software, the system can easily be scaled to monitor an arbitrary amount of rearing vials simultaneously. We evaluated our system on a curated image dataset comprising nearly 470,000 annotated objects and performed several studies on real world experiments. We successfully reproduced results from well-established circadian experiments by comparing the eclosion periods of wild type flies to the clock mutants pershort, perlong and per0 without involvement of any manual labor. Furthermore we show, that the Hatching-Box is able to extract additional information about group behavior as well as to reconstruct the whole life-cycle of the individual specimens. These results not only demonstrate the applicability of our system for long-term experiments but also indicate its benefits for automated monitoring in the general cultivation process.
@@ 1 Main
For many years, Drosophila melanogaster has been employed as a model organism in a broad range of different research domains including genetics, neurosciences, behavioral, cell biology and toxicological studies but also in medical studies developing novel treatments for cancer or other diseases [1, 2]. This kind of research involves experiments examining the physical condition and well-being of Drosophila throughout their whole life cycle. During this process, researchers survey when and how many animals enter a specific developmental stage and how long they remain in this stage. A prominent example for such experiments emerge in the field of circadian research, established primarily by Konopka and Benzer, who investigated changes in the activity rhythm of Drosophila mutants using their eclosion behavior as an indicator [3]. Additionally to the shift in the eclosion rhythm, these clock mutants also show diverging pace in their development as a whole [4].
Despite the growing necessity in various research domains, circadian experiments continue to be a challenging and time-consuming task to this day since the tools available either require fluorescent dyes to visually mark interesting brain regions [5] or extracting individual animals from their rearing vials and monitoring them in a distinctive system. An example of such a system is the commonly used Trikinetics Eclosion Monitor for which Drosophila pupae are glued to a disc and emerging adult flies fall down and are counted as they cross an infrared barrier [6]. Other, camera-based systems, record videos of the animals which can then be annotated semi-automatically by the experimenter with an imaging software such as Fiji [7, 8, 9]. For more detailed studies automated tracking systems have been developed which automatically extract the behavior of individual flies or larvae over time and were surveyed by Panadeiro et al. [10]. For example anTrax, Ctrax and Idtracker.ai are commonly used tracking applications that allow tracking of multiple Drosophila over a period of hours [11, 12, 13]. Providing a fine-tuned combination of a custom made imaging system and tracking software, Risse et al. proposed FIMTrack [14]. The proposed imaging hardware consists of an arena made from an acrylic glass plate which is illuminated by infrared LEDs using frustrated total internal reflection (FTIR) and recorded with an infrared sensitive camera. This system was extended to be used for Drosophila larvae crawling in FTIR-illuminated rearing vials, yielding fine-grained 3-dimensional trajectories of the animals’ movement [15]. Similarly for the domain of ethology Geissmann et al. designed an open source ethoscope system which can track adult flies [16] and analyze their behavior. Apart from methods focused on Drosophila there is also DeepLabCut which is often used for mice [17] but is also applicable for other animals including flies.
@@
Another substantial constraint of these methods with regards to long-term behavior experiments is their inability to discriminate between different developmental stages of Drosophila which is a prerequisite to study the life cycle of flies.
Figure 1: System overview showing stack-able Hatching-Boxes. rearing vials with Drosophila can be placed in the Hatching-Box which then automatically tracks the populations’ behavior and provides images and behavior analysis to a central computer.
To enable high-throughput, long-term (circadian) behavior experiments as well as automatic monitoring of Drosophila development during breeding, we propose an open-source system called Hatching-Box. Our proposed system provides numerous advantages compared to current state-of-the-art: 1. Life-time monitoring of Drosophila is facilitated in off-the-shelf rearing vials during the conventional rearing routine. As a consequence, our system does not require the manual collection of animals and no preparation of the tracking system is necessary, enabling behavioral quantifications without any labor overhead. 2. Each system houses up to three standard (41.5mm diameter) rearing vials, can be placed in incubators or cultivation rooms and only off-the-shelf hardware components were used so that an arbitrary amount of stackable monitoring systems can run in parallel. 3. The self-contained hardware design renders the internal imaging modalities independent from external influences. Hence, the hardware, firmware and software are precisely aligned to fit this setup to yield accurate and reproducible results. 4. A machine learning-based object detection algorithm based on YOLO is adapted and trained on nearly 470.000 manually labeled objects. Our system is capable of accurately detecting all developmental stages of Drosophila even in very cluttered conditions, yielding an accuracy of up to 91% while providing real-time processing capabilities. 5. An array of built-in sensors is integrated to monitor the temperature and humidity within each box. In addition, dedicated light stimulation has been integrated, which can either automatically adapt the inner lightning conditions to the external brightness or can be programmed to provide individual illumination schemes for each box. 6. We implemented an easy-to-use GUI that can be used for controlling the camera, start or stop a recording and automatically analyze the data. 7. All components (i.e. hardware, firmware, software) have been tested and a rigorous evaluation of our system was conducted by reproducing the results of Konopka and Benzer in a well-established circadian experiment by using the three Drosophila clock mutants pershort, perlong and per0 and comparing their activity cycles with the wild type.
@@ 2.1 Hardware
Figure 2: Overview of the proposed hardware architecture. a, A single Hatching-Box consists of a RaspberryPi 4 with the HQ camera module, an LED controller device (Arduino Uno + custom shield), an Arduino Sense BLE 33 sensor board, as well as a light guide panel, outfitted with IR and white LEDs. b, Image as taken by our Hatching-Box with different objects highlighted: third instar larva (cyan), adult fly (blue), empty pupa (red), out-of-focus (yellow) and full pupa (green).
@@
The hardware of our system is designed to be easily integrated into established daily laboratory routines to provide long-term monitoring capabilities without the need for additional manual labor. This requires the use of standard rearing vials for animal housing and a small overall footprint of the system to fit into standard incubators. Additional sensors and light sources provide supervision and control over the breeding conditions in each box. We use a standard small-sized plastic electrical box for the main compartment, housing our central computing device, a Raspberry Pi 4, a camera and additional necessary technical components and up to three standard rearing vials (see Figure 2(a)). The form factor of these boxes enable arbitrary stacking of multiple setups for high-throughput experiments and each box is individually connected via Ethernet to operate several systems in parallel (see Figure 1). The buildtn camera captures frames using a user-specified frame rate allowing for continuous recordings or time-lapse sequences. For illumination an acrylic glass-based light guide panel is positioned behind the rearing vials which is equipped with several near infrared and white light-emitting diodes (LEDs). In order to visually detect small and semi-translucent objects (e.g. empty pupae, larvae, etc.) we optimized the transmitted light configuration by deriving a custom diffusion pattern, providing homogeneous diffuse illumination across the panel (see Appendix C). The IR LEDs are used for image capturing only given this wavelength is reported to be invisible for flies [18]. We synchronized the IR LED activation with the camera so that IR light is only present during image acquisition while being turned off otherwise to prevent heat buildup and to keep a low energy profile (see Sec. 4.1.3). Temperature and humidity sensors inside each box record temporally synchronized measures of the breeding conditions. The white LEDs provide a visual stimuli and can either be freely programmed for each box individually or can be synchronized with the outside brightness using an externally mounted brightness sensor (e.g. to mimic the lightning conditions of the surrounding incubator; see Sec. 4.1.3). The combination of the high resolution camera and the optimized backlight panel yield detailed images of all developmental stages of Drosophila (see Figure 2(b)) which are subsequently used for data collection and in our classification and tracking pipeline.
@@ 2.2 Dataset
For training our YOLOv7 object detector we curated a dataset containing Drosophila in various developmental stages, i.e. larva, pupa, empty pupa, adult fly. In addition we marked out-of-focus objects on the backside of the vials since these blurred objects can be the cause of misdetections and misclassifications. In total we annotated 469,120 objects (bounding boxes and class labels) in 1505 images recorded in different Hatching-Box systems. The first 200 images were labelled thoroughly by domain experts. We utilized these first images to train a first YOLOv7-E6E model which was subsequently used to provide suggestions to the annotators to speed up the labeling process. Importantly, all suggestions proposed by our model were manually checked by domain experts. Table 3(b) shows the class distribution across our dataset.
@@ 2.3.3 Performance evaluation
The experiments were performed in a temperature controlled incubator (CLF, Plant Climatics) set at 25°C, in constant darkness to be able to study the inner rhythm of the flies without effect of the light. We monitored three vials of each genotype per Hatching-Box with a image capturing interval of 10 minutes and configured the tracker to only track full and empty pupae. With this configuration we monitored the different genotypes for 14 days each and computed the quantitative results as described in Sec. 2.3.2. To find the time points of eclosion/pupation based on this output we processed the sequence of detected developmental stages for each separate specimen with a sliding window of size $\tau$ (with $\tau=7$ for our framerate). We mark a time point $t$ as time of eclosion/pupation if a majority of time points $\left[t,\ldots,t+\tau-1\right]$ are associated with the class of empty/full pupa.
Figure 4: 4(a), Positions of pupations (wildtype Drosophila) over the course of our 14 days circadian experiment. 4(b), Larva locomotion (wildtype Drosophila) over 100 second captured with our system with 1 fps.
@@ 3 Discussion
The software is developed with a particular focus on throughput to detect, classify and track specimens in the captured images over the whole course of the experiment without manual intervention. The combination of these different components result in high quality tracking results for experiments with hundreds of animals spanning multiple weeks that are not straightforward to acquire with software-only tracking solutions (e.g. [13, 17]) which require the experimenter to establish an appropriate imaging setup. Being able to track Drosophila in their rearing vials takes the burden off of experimenters to individualize their specimen and simultaneously makes the Hatching-Box also applicable for monitoring of their general development in cultivation rooms and gathering behavioral features on-the-fly.
@@ 4.1.1 Hardware Design
As the main compartment for our system we used an off-the-shelf electrical box measuring 355.0mm ×248.0mm ×110.7mm (length ×width ×height), providing enough room for the necessary electrical components as well as three standard rearing vials with a diameter of 41.5mm (see Appendix B). For our system we employed an adapted set of imaging components, backlight illumination and additional sensors as specified below.
@@ 4.1.2 Imaging components
Image recording is performed with the Raspberry Pi High Quality Camera module connected to the Raspberry Pi 4 (8 GB). The camera module uses the 12.3 megapixel Sony IMX477 sensor and is equipped with a C/CS mount which provides compatibility with a large number of camera lenses. When the camera is used with its full resolution of 4056 ×3040 pixels, our system can achieve a frame rate of up to 1 frame per second. For our experiments we used the C125-0818-5M Basler lens with 8mm focal length due to its minimal working distance (MOD) of 100mm. We outfitted the lens with a MIDOPT LP818-46 low-pass IR filter with a useful range of 825nm - 1100nm. The builtin IR block filter of the camera module was removed.
@@ 4.2.2 Client architecture: HB-client
In our proposed architecture, the term “client” corresponds to an individual Hatching-Box (see Sec. 4.1). The soft- and hardware of each Hatching-Box is controlled by a Raspberry Pi 4. The Raspberry Pi is connected to a Raspberry Pi HQ camera module and, via USB, to an Arduino Sense BLE 33 sensor board which are used for capturing images and sensor data respectively. As a driver for the camera, the HB-client uses the libcamera software stack [27]. Libcamera enables control of the camera’s recording parameters, e.g. exposure time and analog/digital gain, and construction of an image processing pipeline that uses firmware-based image processing in combination with custom-made processing routines. This processing includes routines for isolating a region of interest (ROI) from the captured image as well as detecting the boundaries of each monitored vial to support experimental designs which require observing different classes such as varying phenotypes. The parameters of the processing steps can be controlled in the HB-server application (see Sec. 4.2.3). Images are recorded using the BGR888 pixel format, encoded and compressed in PNG format and subsequently sent to the server. The LED controller, also connected via USB, is specifically designed for circadian experiments and operates the backlight panel. The connected LED strips with different wavelengths (here: white and IR LEDs) can be controlled separately. The HB-client is launched automatically upon startup of the Raspberry Pi and first initializes the serial communication channels to the sensor board, the LED controller and the libcamera software-stack. After initialization, the HB-client connects to the configured IP address and port of the HB-server. Afterwards, the Hatching-Box maintains the connection to the server, reconnects if necessary, and idles until the user issues a snapshot/recording task, changes camera settings or until the next scheduled video frame has to be captured and transmitted.

### v3

# The Hatching-Box: A Novel System for Automated Monitoring and Quantification of Drosophila melanogaster Developmental Behavior

@@ Abstract
Drosophila melanogaster is among the most frequently used model organism in biology and life science as well as for development of novel treatments in the medical domain. In these fields, an analysis of behavioral patterns of the flies throughout their whole life cycle is of significant importance. For experimental purposes, large numbers of different Drosophila genotypes are cultivated in rearing vials and monitored for anomalies in their behavior or development, for example when exposed to stimuli or substances. From this, researchers are able to draw conclusions regarding the effect of the involved substances or treatments on humans, other animals or the environment. A major drawback of current state-of-the-art monitoring systems is the high amount of manual and time-consuming labor as most of the methods employed today require individualization of the flies or similar steps which prevents fully automated experiments. In this paper we propose the Hatching-Box, a novel imaging and analysis system to automatically monitor and quantify the developmental behavior of Drosophila in standard rearing vials and during regular rearing routines, rendering explicit experiments obsolete. This is achieved by combining custom tailored imaging hardware with dedicated detection and tracking algorithms, enabling the quantification of larvae, filled/empty pupae and flies over multiple days. Given the affordable and reproducible design of the Hatching-Box in combination with our generic client/server-based software, the system can easily be scaled to monitor an arbitrary amount of rearing vials simultaneously. We evaluated our system on a curated image dataset comprising nearly 470,000 annotated objects and performed several studies on real world experiments. We successfully reproduced results from well-established circadian experiments by comparing the eclosion periods of wild type flies to the clock mutants pershort, perlong and per0 without involvement of any manual labor. Furthermore we show, that the Hatching-Box is able to extract additional information about group behavior as well as to reconstruct the whole life-cycle of the individual specimens. These results not only demonstrate the applicability of our system for long-term experiments but also indicate its benefits for automated monitoring in the general cultivation process.
@@ 1 Main
The model organism Drosophila melanogaster has been used to study different aspects of biology, such as genetics, neuroscience or cell biology, and in recent years even for translational studies for human diseases [1, 2]. The whole life cycle of Drosophila comprises 4 different stages: egg, larvae, pupae and adults, and often it is of interest to know when and how many animals enter a specific developmental stage and how long they do remain in this stage. One commonly studied transition is the adult emergence or eclosion of Drosophila. In 1971, Konopka and Benzer found that this is time of day dependent and that mutations in a single gene named period affected this process [3]. The mutants they used do not only show a change on the timing of eclosion, but also show diverging pace in their development as a whole [4].
The study of eclosion is still a common procedure to study the circadian rhythms of the flies [5], as well as the developmental timing of pupation [6]. This type of experiments are time-consuming and often involve the constant personal monitoring of the rearing vials. Despite the growing necessity in various research domains, circadian experiments continue to be a challenging and time-consuming task to this day since the tools available either require fluorescent dyes to visually mark interesting brain regions [7] or extracting individual animals from their rearing vials and monitoring them in a distinctive system. An example of such a system is the commonly used Trikinetics Eclosion Monitor for which Drosophila pupae are glued to a disc and emerging adult flies fall down and are counted as they cross an infrared barrier [8]. Other, camera-based systems, record videos of the animals which can then be annotated semi-automatically by the experimenter with an imaging software such as Fiji [9, 10, 11]. For more detailed studies automated tracking systems have been developed which automatically extract the behavior of individual flies or larvae over time and were surveyed by Panadeiro et al. [12]. For example anTrax, Ctrax and Idtracker.ai are commonly used tracking applications that allow tracking of multiple Drosophila over a period of hours [13, 14, 15]. Providing a fine-tuned combination of a custom made imaging system and tracking software, Risse et al. proposed FIMTrack [16]. The proposed imaging hardware consists of an arena made from an acrylic glass plate which is illuminated by infrared LEDs using frustrated total internal reflection (FTIR) and recorded with an infrared sensitive camera. This system was extended to be used for Drosophila larvae crawling in FTIR-illuminated rearing vials, yielding fine-grained 3-dimensional trajectories of the animals’ movement [17]. Similarly for the domain of ethology Geissmann et al. designed an open source ethoscope system which can track adult flies [18] and analyze their behavior. Apart from methods focused on Drosophila there is also DeepLabCut which is often used for mice [19] but is also applicable for other animals including flies.
@@
Another substantial constraint of these methods with regards to long-term behavior experiments is their inability to discriminate between different developmental stages of Drosophila which is a prerequisite to study the life cycle of flies. Since 2023, flyGear offers a commercially available system that can be used to classify Drosophila larvae and pupae, counts them and provides automatic data analysis and visualization.
Figure 1: System overview showing stackable Hatching-Boxes. Standardized rearing vials used to house Drosophila can be placed in the Hatching-Box, which then automatically tracks the population’s behavior and provides images and behavior analysis to a central computer.
To enable high-throughput, long-term (circadian) behavior experiments as well as automatic monitoring of Drosophila development during breeding, we propose an open-source system called Hatching-Box. Our system provides numerous advantages compared to current state-of-the-art: 1. Life-time monitoring of Drosophila is facilitated in off-the-shelf rearing vials during the conventional rearing routine. As a consequence, our system does not require the manual collection of animals and no preparation of the tracking system is necessary, enabling behavioral quantifications without any labor overhead. 2. Each system houses up to three rearing vials (41.5mm diameter), can be placed in incubators or cultivation rooms and comprises primarily off-the-shelf hardware components with some 3D printed and manufactured parts from easily obtainable materials so that an arbitrary amount of stackable monitoring systems can run in parallel. 3. The self-contained hardware design renders the internal imaging modalities independent from external influences. Hence, the hardware, firmware and software are precisely aligned to fit this setup to yield accurate and reproducible results. 4. A machine learning-based object detection algorithm based on YOLO is adapted and trained on nearly 470.000 manually labeled objects. Our system is capable of accurately detecting all developmental stages of Drosophila even in very cluttered conditions, yielding an accuracy of up to 91% while providing real-time processing capabilities. 5. An array of built-in sensors is integrated to monitor the temperature and humidity within each box. In addition, dedicated light stimulation has been integrated, which can either automatically adapt the inner lightning conditions to the external brightness or can be programmed to provide individual illumination schemes for each box. 6. We implemented an easy-to-use GUI that can be used for controlling the camera, start or stop a recording and automatically analyze the data. 7. All components (i.e. hardware, firmware, software) have been tested and a rigorous evaluation of our system was conducted by reproducing the results of Konopka and Benzer in a well-established circadian experiment by using the three Drosophila clock mutants pershort, perlong and per0 and comparing their activity cycles with the wild type. 8. The full source code, collected dataset, a conclusive bill of materials and required files for the 3D printed and manufactured parts will be made available under an open-source license.
@@ 2.1 Hardware
Figure 2: Overview of the proposed hardware architecture. 2(a) A single Hatching-Box consists of a RaspberryPi 4 with the HQ camera module, an LED controller device (Arduino Uno + custom shield), an Arduino Sense BLE 33 sensor board, as well as a light guide panel, outfitted with IR and white LEDs. 2(b) Picture of a Hatching-Box with cover removed. 2(c) Image as taken by our Hatching-Box with different objects highlighted: third instar larva (cyan), adult fly (blue), empty pupa (red), out-of-focus (yellow) and full pupa (green).
@@
The hardware of our system is designed to be easily integrated into established daily laboratory routines to provide long-term monitoring capabilities without the need for additional manual labor. This requires the use of rearing vials for animal housing and a small overall footprint of the system to fit into most incubators. While we use rearing vials with a 41.5mm diameter, our system can be easily adapted to other vials as well by editing one of the provided parametric manufacturing files (see Appendix B). Additional sensors and light sources provide supervision and control over the breeding conditions in each box. We use a standard small-sized plastic electrical box for the main compartment, housing our central computing device, a Raspberry Pi 4, a camera and additional necessary technical components and up to three standard rearing vials (see Figure 2(a)). The form factor of these boxes enable arbitrary stacking of multiple setups for high-throughput experiments and each box is individually connected via Ethernet to operate several systems in parallel (see Figure 1). The built in camera captures frames using a user-specified frame rate allowing for continuous recordings or time-lapse sequences. For illumination an acrylic based light guide panel is positioned behind the rearing vials which is equipped with several near infrared and white light-emitting diodes (LEDs). In order to visually detect small and semi-translucent objects (e.g. empty pupae, larvae, etc.) we optimized the transmitted light configuration by deriving a custom diffusion pattern, providing homogeneous diffuse illumination across the light guide panel (see Appendix D). The IR LEDs are used for image capturing only given this wavelength is reported to be invisible for flies [20]. An additional IR shield in the middle of the compartment prevents light emitted by the computing hardware to have an influence on the specimens or produce reflections on the rearing vials. We synchronized the IR LED activation with the camera so that IR light is only present during image acquisition while being turned off otherwise to prevent heat buildup and to keep a low energy profile (see Sec. 4.1.3). Temperature and humidity sensors inside each box record temporally synchronized measures of the breeding conditions. The white LEDs provide a visual stimuli and can either be freely programmed for each box individually or can be synchronized with the outside brightness using an externally mounted brightness sensor (e.g. to mimic the lightning conditions of the surrounding incubator; see Sec. 4.1.3). The combination of the high resolution camera and the optimized backlight panel yield detailed images of all developmental stages of Drosophila (see Figure 2(c)) which are subsequently used for data collection and in our classification and tracking pipeline.
@@ 2.2 Dataset
For training our YOLOv7 object detector we curated a dataset containing Drosophila in various developmental stages, i.e. larva, pupa, empty pupa, adult fly. In addition we marked out-of-focus objects on the backside of the vials since these blurred objects can be the cause of misdetections and misclassifications. In total we annotated 469,120 objects (bounding boxes and class labels) in 1505 images recorded in different Hatching-Box systems. The first 200 images were labelled thoroughly by domain experts. We utilized these first images to train an initial YOLOv7-E6E model which was subsequently used to provide suggestions to the annotators to speed up the labeling process. Importantly, all suggestions proposed by our model were manually checked by domain experts. Table 3(b) shows the class distribution across our dataset.
@@ 2.3.3 Performance evaluation
The experiments were performed in a temperature controlled incubator (CLF, Plant Climatics) set at 25°C, in constant darkness to be able to study the inner rhythm of the flies without effect of the light. We monitored three vials of each genotype per Hatching-Box with a image capturing interval of 10 minutes and configured the tracker to only track full and empty pupae. With this configuration we monitored the different genotypes for 14 days each and computed the quantitative results as described in Sec. 2.3.2. To find the time points of eclosion/pupation based on this output we processed the sequence of detected developmental stages for each separate specimen with a sliding window of size $\tau$ (with $\tau=7$ for our framerate). We mark a time point $t$ as time of eclosion/pupation if a majority of time points $\left[t,\ldots,t+\tau-1\right]$ are associated with the class of empty/full pupa. For the performed circadian experiments we used this approach to extract the eclosion time points only.
Figure 4: 4(a), Positions of eclosion (wildtype Drosophila) over the course of our 14 days circadian experiment. 4(b), Larva locomotion (wildtype Drosophila) over 100 second captured with our system with 1 fps.
@@ 3 Discussion
The software is developed with a particular focus on throughput to detect, classify and track specimens in the captured images over the whole course of the experiment without manual intervention. The combination of these different components result in high quality tracking results for experiments with hundreds of animals spanning multiple weeks that are not straightforward to acquire with software-only tracking solutions (e.g. [15, 19]) which require the experimenter to establish an appropriate imaging setup. The presence of user-controllable lights within each Hatching-Box facilitates multiple experiments with different light cycles in the same incubator. Additionally, by utilizing our sensor array, the environmental conditions during the experiment can be closely monitored and easily correlated with the obtained developmental data. Being able to track Drosophila in their rearing vials takes the burden off of experimenters to individualize their specimen and simultaneously makes the Hatching-Box also applicable for monitoring of their general development and gathering behavioral and fitness features on-the-fly.
@@ 4.1.1 Hardware Design
As the main compartment for our system we used an off-the-shelf electrical box measuring 355.0mm ×248.0mm ×110.7mm (length ×width ×height), providing enough room for the necessary electrical components as well as three rearing vials with a diameter of 41.5mm (see Appendix C). For our system we employed an adapted set of imaging components, backlight illumination and additional sensors as specified below (see Appendix B).
@@ 4.1.2 Imaging components
Image recording is performed with the Raspberry Pi High Quality Camera module connected to the Raspberry Pi 4 (8 GB). The camera module uses the 12.3 megapixel Sony IMX477 sensor and is equipped with a C/CS mount which provides compatibility with a large number of camera lenses. When the camera is used with its full resolution of 4056 ×3040 pixels, our system can achieve a frame rate of up to 1 frame per second. For our experiments we used the C125-0818-5M Basler lens with 8mm focal length due to its minimal working distance (MOD) of 100mm. We outfitted the lens with a MIDOPT LP818-46 low-pass IR filter with a useful range of 825nm - 1100nm. The builtin IR block filter of the camera module was removed. Additionally
@@ 4.2.2 Client architecture: HB-client
In our proposed architecture, the term “client” corresponds to an individual Hatching-Box (see Sec. 4.1). The soft- and hardware of each Hatching-Box is controlled by a Raspberry Pi 4. The Raspberry Pi is connected to a Raspberry Pi HQ camera module and, via USB, to an Arduino Sense BLE 33 sensor board which are used for capturing images and sensor data respectively. As a driver for the camera, the HB-client uses the libcamera software stack [29]. Libcamera enables control of the camera’s recording parameters, e.g. exposure time and analog/digital gain, and construction of an image processing pipeline that uses firmware-based image processing in combination with custom-made processing routines. This processing includes routines for isolating a region of interest (ROI) from the captured image as well as detecting the boundaries of each monitored vial to support experimental designs which require observing different classes such as varying phenotypes. The parameters of the processing steps can be controlled in the HB-server application (see Sec. 4.2.3). Images are recorded using the BGR888 pixel format, encoded and compressed in PNG format and subsequently sent to the server. The LED controller, also connected via USB, is specifically designed for circadian experiments and operates the backlight panel. The connected LED strips with different wavelengths (here: white and IR LEDs) can be controlled separately. The HB-client is launched automatically upon startup of the Raspberry Pi and first initializes the serial communication channels to the sensor board, the LED controller and the libcamera software-stack. After initialization, the HB-client connects to the configured IP address and port of the HB-server. Afterwards, the Hatching-Box maintains the connection to the server, reconnects if necessary, and idles until the user issues a snapshot/recording task, changes camera settings or until the next scheduled video frame has to be captured and transmitted. By transferring the recorded images directly to the server, we are able to fully omit write operations to the micro SD card, preventing data corruption.